# Easi3R: Estimating Disentangled Motion from DUSt3R Without Training

### v1

@@ 3.1 DUSt3R with Dynamic Video
Figure 3: DUSt3R and our Easi3R adaptation. DUSt3R encodes two images $I^{a},I^{b}$ into feature tokens $\mathbf{F}_{0}^{a},\mathbf{F}_{0}^{b}$, which are then decoded into point maps in the reference view coordinate space using two decoders. Our Easi3R aggregates the cross-attention maps from the decoders, producing four semantically meaningful maps: $\mathbf{A}^{b=\text{src}}_{\mu},\mathbf{A}^{b=\text{src}}_{\sigma},\mathbf{A}^{a=\text{ref}}_{\mu},\mathbf{A}^{b=\text{ref}}_{\sigma}$. These maps are then used for a second inference pass to enhance reconstruction quality.
@@ 3.2 Secrets Behind DUSt3R
We now examine the network architecture to identify the components that cause failures for dynamic video input. As shown in Figure 3, DUSt3R consists of two branches: the top one for the reference image $I^{a}$ and the bottom one for the source image $I^{b}$. The two input images are first processed by a weight-sharing ViT encoder [11], producing token representations $\mathbf{F}_{0}^{a}=\text{Encoder}(I^{a})$ and $\mathbf{F}_{0}^{b}=\text{Encoder}(I^{b})$. Next, two decoders, composed of a sequence of decoder blocks, exchange information both within and between views. In each block, self-attention is applied to the token outputs from the previous block, while cross-attention is performed using the corresponding block outputs from the other branch,
@@
Spatial attention maps. As illustrated in the Figure 3 (left), the image features $\mathbf{F}$ are projected into a query matrix for their respective branch with $\mathbf{Q}=\ell_{\mathbf{Q}}(\mathbf{F})\in\mathbb{R}^{(h\times w)\times c}$, while also serving as a key and value matrix for the other matrices, $\mathbf{K}=\ell_{\mathbf{K}}(\mathbf{F})\in\mathbb{R}^{(h\times w)\times c}$, where $c$ is the feature dimenson. The projectios are obtained using trainable linear functions $\ell_{\mathbf{Q}}(\cdot),\ell_{\mathbf{K}}(\cdot)$. As illumistracted in the right side of Figure 3, this results in the cross-attention map:
@@ Appendix D Addtional Results
We report additional qualitative results of disentangled 4D reconstruction in Figure 10, Figure 11 and Figure 12. We find that MonST3R tends to predict under-segmented dynamic masks, while DAS3R tends to predict over-segmented dynamic masks. CUT3R, although it produces more accurate depth estimation, is prone to being affected by dynamic objects, leading to misaligned static structures, unstable camera pose estimation, and ghosting artifacts due to the lack of dynamic segmentation prediction. In contrast, Easi3R achieves more accurate segmentation, camera pose estimation, and 4D reconstruction, resulting in renderings with better visual quality.

### v2

@@ 3.1 DUSt3R with Dynamic Video
Figure 3: DUSt3R and our Easi3R adaptation. DUSt3R encodes two images $I^{a},I^{b}$ into feature tokens $\mathbf{F}_{0}^{a},\mathbf{F}_{0}^{b}$, which are then decoded into point maps in the reference view coordinate space using two decoders. Our Easi3R aggregates the cross-attention maps from the decoders, producing four semantically meaningful maps: $\mathbf{A}^{b=\text{src}}_{\mu},\mathbf{A}^{b=\text{src}}_{\sigma},\mathbf{A}^{a=\text{ref}}_{\mu},\mathbf{A}^{a=\text{ref}}_{\sigma}$. These maps are then used for a second inference pass to enhance reconstruction quality.
@@ 3.2 Secrets Behind DUSt3R
We now examine the network architecture to identify the components that cause failures for dynamic video input. As shown in Figure 3, DUSt3R consists of two branches: the top one for the reference image $I^{a}$ and the bottom one for the source image $I^{b}$. The two input images are first processed by a weight-sharing ViT encoder [11], producing token representations $\mathbf{F}_{0}^{a}=\text{Encoder}(I^{a})$ and $\mathbf{F}_{0}^{b}=\text{Encoder}(I^{b})$. Next, two decoders, composed of a sequence of decoder blocks, exchange information both within and between views. In each block, self-attention is applied to the token outputs from the previous block, while cross-attention is performed using the corresponding block outputs from the other branch.
@@
Spatial attention maps. As illustrated in the Figure 3 (left), the image features $\mathbf{F}$ are projected into a query matrix for their respective branch with $\mathbf{Q}=\ell_{\mathbf{Q}}(\mathbf{F})\in\mathbb{R}^{(h\times w)\times c}$, while also serving as a key and value matrix for the other matrices, $\mathbf{K}=\ell_{\mathbf{K}}(\mathbf{F})\in\mathbb{R}^{(h\times w)\times c}$, where $c$ is the feature dimenson. The projections are obtained using trainable linear functions $\ell_{\mathbf{Q}}(\cdot),\ell_{\mathbf{K}}(\cdot)$. As illumistracted in the right side of Figure 3, this results in the cross-attention map:
@@ Appendix C Limitations
Figure 10: Dynamic masks in static scene. Easi3R tends to reweight low confident regions in static scenes, leading to better pose estimation, as shown in Tab. 9.
Sintel BONN KITTI
Alignment Method Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$ <$1.25\uparrow$
Per-sequence scale & shift Marigold 0.532 51.5 0.091 93.1 0.149 79.6
DepthAnythingV2 0.367 55.4 0.106 92.1 0.140 80.4
NVDS 0.408 48.3 0.167 76.6 0.253 58.8
ChronoDepth 0.687 48.6 0.100 91.1 0.167 75.9
DepthCrafter 0.292 69.7 0.075 97.1 0.110 88.1
Robust-CVD 0.703 47.8 - - - -
CasualSAM 0.387 54.7 0.169 73.7 0.246 62.2
MASt3R 0.327 59.4 0.167 78.5 0.137 83.6
Spann3R 0.508 50.8 0.157 82.1 0.207 73.0
CUT3R 0.454 55.7 0.074 94.5 0.106 88.7
DUSt3R 0.531 51.2 0.156 83.1 0.135 81.8
MonST3R 0.333 59.0 0.066 96.4 0.157 73.8
[…]
Per-sequence scale MASt3R 0.641 43.9 0.252 70.1 0.183 74.5
Spann3R 0.622 42.6 0.144 81.3 0.198 73.7
CUT3R 0.421 47.9 0.078 93.7 0.118 88.1
DUSt3R 0.656 45.2 0.155 83.3 0.144 81.3
MonST3R 0.378 55.8 0.067 96.3 0.168 74.4
[…]
Table 8: Video Depth Evaluation. We use the evaluation results from CUT3R for baselines.
Sintel TUM-dynamics ScanNet (static)
Category Method ATE $\downarrow$ RPE trans $\downarrow$ RPE rot $\downarrow$ ATE $\downarrow$ RPE trans $\downarrow$ RPE rot $\downarrow$ ATE $\downarrow$ RPE trans $\downarrow$ RPE rot $\downarrow$
Pose only DROID-SLAM 0.175 0.084 1.912 - - - - - -
DPVO 0.115 0.072 1.975 - - - - - -
Particle-SfM 0.129 0.031 0.535 - - - 0.136 0.023 0.836
LEAP-VO 0.089 0.066 1.250 0.068 0.008 1.686 0.070 0.018 0.535
Pose & Depth Robust-CVD 0.360 0.154 3.443 0.153 0.026 3.528 0.227 0.064 7.374
CasualSAM 0.141 0.035 0.615 0.071 0.010 1.712 0.158 0.034 1.618
MASt3R 0.185 0.060 1.496 0.038 0.012 0.448 0.078 0.020 0.475
Spann3R 0.329 0.110 4.471 0.056 0.021 0.591 0.096 0.023 0.661
CUT3R 0.213 0.066 0.621 0.046 0.015 0.473 0.099 0.022 0.600
DUSt3R 0.417 0.250 5.796 0.083 0.017 3.567 0.081 0.028 0.784
MonST3R 0.111 0.044 0.869 0.098 0.019 0.935 0.077 0.018 0.529
[…]
Table 9: Camera Pose Evaluation. We use the evaluation results from CUT3R for baselines.
Input Output DAVIS-16 DAVIS-17
[…]
Supervised SFL 67.4 - - -
SIMO 67.8 - - -
OCLR-flow 72.0 70.0 - 69.9
OCLR-TTA 80.8 80.2 - 76.0
FlowSAM 87.1 85.7 - -
SegAnyMo 90.6 89.2 - 90.0
DAS3R 54.2 51.6 57.4 55.5
Unsupervised SAGE 42.6 - - -
CUT 55.2 - - -
FTS 55.8 - - -
CIS 70.3 67.6 - -
Motion Grouping 68.3 - - -
EM 69.3 76.2 - -
RCF-Stage1 80.2 78.6 - -
RCF-All 82.1 81.0 - -
LRTL 82.2 - - -
Zero-shot MonST3R 64.3 61.4 56.4 59.0
[…]
Table 10: Comparisons of Dynamic Object Segmentation on DAVIS with 2D dynamic segmentation methods.
DAVIS-16 DAVIS-17 DAVIS-all
[…]
Window Size 3 76.0 89.2 70.8 82.4 65.7 76.2
5* 70.7 79.9 67.9 76.1 63.1 72.6
7 66.9 76.9 63.9 73.3 61.0 68.8
Number of Clusters 16 67.4 79.1 64.6 73.7 60.7 65.2
32 71.6 83.9 68.0 78.3 65.1 75.2
64* 70.7 79.9 67.9 76.1 63.1 72.6
128 69.9 79.9 66.3 76.2 62.9 73.1
Thresholding Values 0.5 61.6 64.5 61.0 65.2 61.6 67.8
0.7 70.2 85.0 62.8 71.8 58.1 67.0
Otsu’s method* 70.7 79.9 67.9 76.1 63.1 72.6
Table 11: More ablations on segmentation quality using DAVIS. * denotes the value used in the submission.
@@ Appendix D Addtional Results
More Evaluation. We evaluated on full-length sequences with downsampling only for GPU fit, leading to more dynamic and challenging motion than prior short-clip settings. For a more general evaluation, we also include MonST3R and CUT3R evaluation protocols. Tab. 8 and Tab. 9 confirm the effectiveness.
Behavior for Static Scenes. Interestingly, our method also improves on static scenes (ScanNet in Tab. 9 and Fig. 10), owing to our attention reweighting.
Comparison with 2D Baselines. We also include a comparison with 2D baselines. As shown in Tab. 10, Easi3R achieves SOTA segmentation in a zero-shot manner with only the image as input.
More Ablation. We further ablate the default settings - window size of 5, 64 clusters, and Otsu’s method in Tab. 11. Using the recent segmentation method SegAnyMo, pose accuracy improves by 9.62% and depth accuracy improves by 4.11% on the Sintel dataset.
Runtime Our method runs at almost the same speed as MonST3R. MonST3R runs at $0.33$ FPS, while Easi3R achieves $0.31$ FPS for $512\times 144$ image resolution on an NVIDIA RTX 4090 GPU.
Qualitative Results We report additional qualitative results of disentangled 4D reconstruction in Figure 11, Figure 12 and Figure 13. We find that MonST3R tends to predict under-segmented dynamic masks, while DAS3R tends to predict over-segmented dynamic masks. CUT3R, although it produces more accurate depth estimation, is prone to being affected by dynamic objects, leading to misaligned static structures, unstable camera pose estimation, and ghosting artifacts due to the lack of dynamic segmentation prediction. In contrast, Easi3R achieves more accurate segmentation, camera pose estimation, and 4D reconstruction, resulting in renderings with better visual quality.